If the measures of the three angles of a triangle are x,2x-20, and 3x-10, then what is this type of triangle?Why?
Answer by mananth(16946)

x+2x-20+ 3x-10= 180
6x-30=180
add 30
6x-30+30=180+30
6x=210
x= 210/6
x= 35
second angle = 2x-20
=2*35-20
=70-20
=50
..
Third angle = 3x-10
3*35-10
105-10
=95
No angles are same. so scalene triangle